# Deterministic Policies for Constrained Reinforcement Learning in Polynomial-Time

### v1

@@ Abstract
We present a novel algorithm that efficiently computes near-optimal deterministic policies for constrained reinforcement learning (CRL) problems. Our approach combines three key ideas: (1) value-demand augmentation, (2) action-space approximate dynamic programming, and (3) time-space rounding. Under mild reward assumptions, our algorithm constitutes a fully polynomial-time approximation scheme (FPTAS) for a diverse class of cost criteria. This class requires that the cost of a policy can be computed recursively over both time and (state) space, which includes classical expectation, almost sure, and anytime constraints. Our work not only provides provably efficient algorithms to address real-world challenges in decision-making but also offers a unifying theory for the efficient computation of constrained deterministic policies.
@@ 1 Introduction
Constrained Reinforcement Learning (CRL) traditionally produces stochastic, expectation-constrained policies that can behave undesirably - imagine a self-driving car that randomly changes lanes or runs out of fuel. However, artificial decision-making systems must be predictable, trustworthy, and robust. One approach to ensuring these qualities is to focus on deterministic policies, which are inherently predictable and trustworthy. Moreover, they are easy to implement [10], reliable for autonomous vehicles [16, 12], and effective for multi-agent coordination [23]. Similarly, almost sure and anytime constraints [21] provide inherent trustworthiness and robustness, essential for applications in medicine [6, 22, 18], disaster relief [9, 29, 27], and resource management [20, 19, 24, 4]. Despite the advantages of deterministic policies and stricter constraints, their computation remains an open challenge in CRL. Our research aims to address this challenge by studying the computational complexity of computing deterministic policies for a wide range of constraint types.
@@
Although optimal stochastic policies for expectation-constrained problems are efficiently computable, the situation drastically changes when we require deterministic policies and general constraints. Computing optimal deterministic policies is NP-hard for most popular constraints, including expectation [10], chance [30], almost sure, and anytime constraints [21]. This complexity remains even if we relax our goal to finding just one feasible policy, provided that we are dealing with a single chance constraint [30], or any two of the other mentioned constraints [21]. Beyond these computational challenges, traditional solution methods, such as backward induction [25, 2], fail to apply due to the cyclic dependencies among subproblems: the value of any decision may depend on the costs of both preceding and concurrent decisions, preventing a solution from being computed in a single backward pass. Therefore, our problem is far from trivial and calls for novel, sophisticated techniques.
@@
Previous work has proposed numerous algorithms, yet none simultaneously achieve computational efficiency, feasibility, and optimality. Optimal and feasible algorithms, albeit inefficient, utilize Mixed-Integer Linear Programs [8] and Dual-guided heuristic forward searches [15] for expectation constraints, and cost-augmented MDPs for almost sure [5] and anytime constraints [21]. Conversely, optimal and efficient algorithms, though infeasible, are known for expectation [26], almost sure, and anytime constraints [21]. A fully polynomial-time approximation scheme (FPTAS) [28] is known for expectation constraints, but it requires strong assumptions such as a constant horizon [17]. Thus, striking a balance between computational efficiency, feasibility, and optimality remains a significant challenge in the pursuit of deterministic policies.
@@
We present an affirmative answer to our question for a class of criteria that we call time-space recursive (TSR), which captures expectation, almost sure, and anytime constraints. We propose a novel algorithm that computes a near-optimal and constraint-compliant deterministic policy for any cMDP and TSR criterion. When the rewards are either polynomially bounded or non-negative, our algorithm becomes an FPTAS, which achieves the best approximation guarantees under worst-case analysis. Our algorithmic approach fully exploits the TSR condition, which enables the cost of a policy to be computed recursively in both time and (state) space. Conversely, several non-TSR criteria, such as chance constraints, are provably inapproximable further highlighting the pivotal role of TSR for efficient computation. Thus, our work not only develops the first provably efficient, near-optimal algorithm for the problem but also sheds light on a fundamental property that enables efficient computation of constrained deterministic policies.
Our algorithmic approach breaks down into three main ideas: (1) value-demand augmentation, (2) action-space approximate dynamic programming, and (3) time-space rounding. We first reduce our problem to an equivalent minimization problem. Then, we augment the states with value demands and the actions with future value demands to break cyclic subproblem dependencies, enabling dynamic programming methods. Importantly, we use values because they can be rounded without compromising feasibility [21] and can capture constraints that are not predictable from realized cumulative costs. However, this results in an exponential action space, making the solution of the Bellman operator as hard as the knapsack problem. By exploiting the space-recursive nature of the criterion, we can efficiently approximate the Bellman operator with dynamic programming. Finally, rounding value demands result in approximation errors over both time and space, but carefully controlling these errors ensures provable guarantees.
@@ 1.1 Related Work
Stochastic packing problems, which generalize the knapsack problem, are often special cases of our problem. Dean et al. [7], Frieze and Clarke [11] derived algorithms with optimal approximation ratio for stochastic and integer packing with multiple constraints, respectively. Yang et al. [31], Bhalgat et al. [3] designed efficient approximation algorithms for variations of the stochastic knapsack problem. Lastly, Halman et al. [14] derived an FPTAS for a general class of stochastic dynamic programs, which was then further improved in [13, 1]. These methods require a single-dimensional state space that captures the constraint. In contrast, our problems have an innate state space in addition to the constraint, making our problem multi-dimensional. Our work can be viewed as a similar general dynamic programming framework but designed for the more complex MDP setting.
@@ 3 Covering Algorithm
The key invariant we desire is that any feasible policy $\pi$ for $\bar{M}$ should satisfy $\bar{V}^{\pi}_{h}(s,v)\geq v$. To ensure this invariant, we recall the policy evaluation equations [25]. If $\pi_{h}(s)=a$, then,
@@ 4 Fast Bellman Updates
To turn this recursion into a usable dynamic programming algorithm, we must also precompute the inputs to any subcomputation. Unlike in standard RL, this computation must be done with a forward recursion. The details for the approximate Bellman update are given in Definition 5.
@@ 5 Approximation Algorithms
Given a rounding down function $\left\lfloor\cdot\right\rfloor_{\mathcal{G}}$, we would ideally use the rounded set $\left\{\left\lfloor v\right\rfloor_{\mathcal{G}}\mid v\in\mathcal{V}\right\}$ to form our approximate state space. To avoid having to compute $\mathcal{V}$ explicitly, we instead use the rounded superset $\left\{\left\lfloor v\right\rfloor_{\mathcal{G}}\mid v\in[v_{min},v_{max}]\right\}$, where $v_{min}$ and $v_{max}$ are bounds on the extremal values that we specify later. To ensure we can use Algorithm 4 to find solutions efficiently, we must also relax the augmented action space to only include vectors that lead to feasible subproblems for (ADP). From Lemma 1, we know this is exactly the set of $(a,\hat{\mathbf{v}})$ for which $\hat{\sigma}_{h,\hat{\mathbf{v}}}^{s,a}(1,r_{h}(s,a))\geq\kappa(v)$. Combining these ideas yields the new approximate MDP, which is defined in Definition 6.
@@
Definition 7 (Additive Rounding).
@@
Definition 8 (Relative Rounding).
@@ 6 Conclusions
In this paper, we studied the computational complexity of computing deterministic policies for CRL problems. Our main contribution was the design of an FPTAS, Algorithm 5, that solves (CON) for any cMPD and TSR criteria under mild reward assumptions. In particular, our method is an additive-FPTAS if the cMDP’s rewards are bounded by a polynomial in the input size, and is a relative-FPTAS if the cMDP’s rewards are non-negative. We note these assumptions are necessary for efficient approximation, so our algorithm achieves the best approximation guarantees possible under worst-case analysis. Moreover, our algorithmic approach, which uses approximate dynamic programming over time and the state space, highlights the importance of the TSR condition in making (CON) tractable. The TSR condition captures the fundamental structure that renders any cost criteria efficiently approximable, thereby serves as a unifying theory for constrained deterministic policies. Thus, our work not only develops the first provably efficient algorithm for (CON) but also uncovers novel algorithmic techniques and structural insights that serve as a foundation for truly practical algorithms.
@@
Several interesting questions remain unanswered. First, it remains unresolved whether an FPTAS asymptotically faster than ours can be constructed. Second, it is unclear whether our TSR condition is necessary for efficient computation or whether a more general condition could be derived. Lastly, it would be desirable to design algorithms that can feasibly handle multiple constraints from Proposition 2. Although computing feasible policies for multiple of these constraints is NP-hard, there could be interesting special cases that are approximable efficiently.

### v2

# Deterministic Policies for Constrained Reinforcement Learning in Polynomial Time

@@ Abstract
We present a novel algorithm that efficiently computes near-optimal deterministic policies for constrained reinforcement learning (CRL) problems. Our approach combines three key ideas: (1) value-demand augmentation, (2) action-space approximate dynamic programming, and (3) time-space rounding. Our algorithm constitutes a fully polynomial-time approximation scheme (FPTAS) for any time-space recursive (TSR) cost criteria. A TSR criteria requires the cost of a policy to be computable recursively over both time and (state) space, which includes classical expectation, almost sure, and anytime constraints. Our work answers three open questions spanning two long-standing lines of research: polynomial-time approximability is possible for 1) anytime-constrained policies, 2) almost-sure-constrained policies, and 3) deterministic expectation-constrained policies.
@@ 1 Introduction
Constrained Reinforcement Learning (CRL) traditionally produces stochastic, expectation-constrained policies that can behave undesirably - imagine a self-driving car that randomly changes lanes or runs out of fuel. However, artificial decision-making systems must be predictable, trustworthy, and robust. One approach to ensuring these qualities is to focus on deterministic policies, which are inherently predictable, easily implemented [19], reliable for autonomous vehicles [30, 23], and effective for multi-agent coordination [38]. Similarly, almost sure and anytime constraints [36] provide inherent trustworthiness and robustness, essential for applications in medicine [15, 37, 32], disaster relief [18, 50, 45], and resource management [35, 34, 40, 7]. Despite the advantages of deterministic policies and stricter constraints, even the computation of approximate solutions has remained an open challenge since NP-hardness was proven nearly 25 years ago  [19]. Our work addresses this challenge by studying the computational complexity of computing deterministic policies for general constraint criteria.
@@
Although optimal stochastic policies for expectation-constrained problems are efficiently computable [3], the situation drastically changes when we require deterministic policies and general constraints. Computing optimal deterministic policies is NP-hard for most popular constraints, including expectation [19], chance [51], almost sure, and anytime constraints [36]. This complexity remains even if we relax our goal to finding just one feasible policy, provided that we are dealing with a single chance constraint [51], or at least two of the other mentioned constraints [36]. Beyond these computational challenges, traditional solution methods, such as backward induction [41, 3], fail to apply due to the cyclic dependencies among subproblems: the value of any decision may depend on the costs of both preceding and concurrent decisions, preventing a solution from being computed in a single backward pass.
@@
Past approaches fail to simultaneously achieve computational efficiency, feasibility, and optimality. Optimal and feasible algorithms, albeit inefficient, utilize Mixed-Integer Linear Programs [17] and Dual-guided heuristic forward searches [29] for expectation-constraints, and cost-augmented MDPs for almost sure [11] and anytime constraints [36]. Conversely, optimal and efficient, though infeasible, algorithms are known for expectation [43], almost sure, and anytime constraints [36]. A fully polynomial-time approximation scheme (FPTAS) [49] is known for expectation constraints, but it requires strong assumptions such as a constant horizon [31]. Balancing computational efficiency, feasibility, and optimality remains the bottleneck to efficient approximation.
@@
We present an FPTAS for computing deterministic policies under any time-space recursive (TSR) constraint criteria. A TSR criteria requires the cost of a policy to be computable recursively in both time and (state) space, which captures expectation, almost sure, and anytime constraints. Since non-TSR criteria, such as chance constraints [51], are provably inapproximable, TSR seems pivotal for efficient computation. Overall, our general framework answers three open complexity questions spanning two longstanding lines of work: we prove polynomial-time approximability for 1) anytime-constrained policies, 2) almost-sure-constrained policies, and 3) deterministic expectation-constrained policies, which have been open for nearly 25 years  [19].
Our approach breaks down into three main ideas: (1) value-demand augmentation, (2) action-space approximate dynamic programming, and (3) time-space rounding. We augment the states with value demands and the actions with future value demands to break cyclic subproblem dependencies, enabling dynamic programming methods. Importantly, we use values because they can be rounded without compromising feasibility [36] and can capture constraints that are not predictable from cumulative costs. However, this results in an exponential action space that makes solving the Bellman operator as hard as the knapsack problem. By exploiting the space-recursive nature of the criterion, we can efficiently approximate the Bellman operator with dynamic programming. Finally, rounding value demands result in approximation errors over both time and space, but carefully controlling these errors ensures provable guarantees.
@@ 1.1 Related Work
Many stochastic packing problems, which generalize the knapsack problem, are captured by our problem. Dean et al. [16], Frieze and Clarke [21] derived optimal approximation ratio algorithms for stochastic packing and integer packing with multiple constraints, respectively. Yang et al. [52], Bhalgat et al. [6] designed efficient approximation algorithms for variations of the stochastic knapsack problem. Then, Halman et al. [27] derived an FPTAS for a general class of stochastic dynamic programs, which was then further improved in [26, 1]. These methods require a single-dimensional state space that captures the constraint. In contrast, our problems have an innate state space in addition to the constraint. Our work forms a similar general dynamic programming framework for the more complex MDP setting.
Constrained RL.
It is known that stochastic expectation-constrained policies are polynomial-time computable via linear programming  [3], and many planning and learning algorithms exist for them  [39, 46, 8, 28]. Some learning algorithms can even avoid violation during the learning process under certain assumptions [48, 4]. Furthermore, Brantley et al. [10] developed no-regret algorithms for cMDPs and extended their algorithms to the setting with a constraint on the cost accumulated over all episodes, which is called a knapsack constraint [10, 13].
Safe RL.
The safe RL community [22, 25] has mainly focused on no-violation learning for stochastic expectation-constrained policies [14, 9, 2, 12, 5] and solving chance constraints [47, 53], which capture the probability of entering unsafe states. Performing learning while avoiding dangerous states [53] is a special case of expectation constraints that has also been studied [42, 44] under non-trivial assumptions. In addition, instantaneous constraints, which require the expected cost to be within budget at all times, have also been studied [33, 20, 24].
@@ 3 Covering Algorithm
The key invariant we desire is that any feasible policy $\pi$ for $\bar{M}$ should satisfy $\bar{V}^{\pi}_{h}(s,v)\geq v$. To ensure this invariance, we recall the policy evaluation equations [41]. If $\pi_{h}(s)=a$, then,
@@ 4 Fast Bellman Updates
To turn this recursion into a usable dynamic programming algorithm, we must also pre-compute the inputs to any sub-computation. Unlike in standard RL, this computation must be done with a forward recursion. The details for the approximate Bellman update are given in Definition 5.
@@ 5 Approximation Algorithms
Given a rounding down function $\left\lfloor\cdot\right\rfloor_{\mathcal{G}}$, we would ideally use the rounded set $\left\{\left\lfloor v\right\rfloor_{\mathcal{G}}\mid v\in\mathcal{V}\right\}$ to form our approximate state space. To avoid having to compute $\mathcal{V}$ explicitly, we instead use the rounded superset $\left\{\left\lfloor v\right\rfloor_{\mathcal{G}}\mid v\in[v_{min},v_{max}]\right\}$, where $v_{min}$ and $v_{max}$ are bounds on the extremal values that we specify later. To ensure we can use Algorithm 4 to find solutions efficiently, we must also relax the augmented action space to only include vectors that lead to feasible subproblems for (ADP). From Lemma 1, we know this is exactly the set of $(a,\hat{\mathbf{v}})$ for which $\hat{\sigma}_{h,\hat{\mathbf{v}}}^{s,a}(1,r_{h}(s,a))\geq\kappa(v)$. Combining these ideas yields the new approximate MDP, defined in Definition 6.
@@
Definition 7 (Additive Approx).
@@
Definition 8 (Relative Approx).
@@ 6 Conclusions
In this paper, we studied the computational complexity of computing deterministic policies for CRL problems. Our main contribution was the design of an FPTAS, Algorithm 5, that solves (CON) for any cMPD and TSR criteria under mild reward assumptions. In particular, our method is an additive-FPTAS if the cMDP’s rewards are polynomially bounded, and is a relative-FPTAS if the cMDP’s rewards are non-negative. We note these assumptions are necessary for efficient approximation, so our algorithm achieves the best approximation guarantees possible under worst-case analysis. Moreover, our algorithmic approach, which uses approximate dynamic programming over time and the state space, highlights the importance of the TSR condition in making (CON) tractable. Our work finally resolves the long-standing open questions of polynomial-time approximability for 1) anytime-constrained policies, 2) almost-sure-constrained policies, and 3) deterministic expectation-constrained policies.
@@
Several interesting questions remain unanswered. First, it remains unresolved whether an FPTAS exists asymptotically faster than ours. Second, whether our TSR condition is necessary for efficient computation or whether a more general condition could be derived is unclear. Lastly, it is open whether there exist algorithms that can feasibly handle multiple constraints from Proposition 2. Although computing feasible policies for multiple constraints is NP-hard, special cases may be approximable efficiently. Moreover, an average-case or smoothed-case analysis could circumvent this worst-case hardness.